OpenStudy (anonymous):

Find 3 consecutive odd integers such that the sum of the first and the third equals the sum of the second and 31.

OpenStudy (anonymous):

x=x. y=x+2, z=x+4... because ODD integers, you skip a number each time. (x+z)=(y+31): (x+(x+4))=((x+2)+31). Solve for x gives you 29. Then you plug in x value for y,z. You get 29,31,33.